Company produces 2 products Π1 and Π2, with a joint function of total cost TC=40Q1+4Q1Q2+40Q2 and a joint function of total revenue TR=200Q1-4Q21+12Q1Q2+120Q2-2Q22, where Q1 and Q2 are the quantities of Π1 and Π2 respectively.

1) Find the maximum profit if the company has a production contract of 60 units in total.

2) Αpproximate the new maximum profit if total output increases by one unit.
Answer #1

TR = 2000,-402 +120,02 +12002 - 207 Page 1 TC = 900, + 40,02 + 4002 Joint porfit (11) - TR - TO a = 2000,- 40,2 +120,02 +1200

Total = 60 Page 2 0,+ 02 - 60 02 = 60 -, - Put ean. in ean. O O, - 0,20 Q, - (60-0,) = 20 Q, -60+Q,- 20 →20, = 80 0,- 40 O2 =

(20.5)^ @,-02-20 Page 3 Q, - (61-0,) = 20 20, = 20 +61 a, = 40.5) Qg = 61-01 = 61-40.5 Q2 = 20.5 at Q, = 40.5 & On = 20.5; TR

**if you liked the answer, then please upvote. Would be motivating for me. Thanks!